Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name V. Sulkosky
Log entry time 09:13:15 on December 22, 2008
Entry number 254390
This entry is a followup to: 254256
keyword=Experiment Summary for Dec 21, Owl - Dec 22, Owl
Summary for Dec 21 owl - Dec 22 owl
Smooth data taking.
Beam in Use
shift |
ABU |
BANU |
BNA |
2008/12/21 owl | 5.14 |
0.13 | 2.74 |
2008/12/21 day | 6.08 |
0.21 | 1.72 |
2008/12/21 swing | 3.88 |
0.55 | 3.57 |
2008/12/22 owl | 6.57 |
0.44 | 0.98 |
total | 21.7 |
1.3 | 9.0 |
Main events/issues:
- Production of pi+ on Vertical pumping at 13.5 uA.
- Dec. 21, 06:00-08:25 MCC vacuum problem.
- Dec. 21, 08:25-09:05 Beam restoration and position check.
- Dec. 21, day shift had over 20 trips per hour until 15:00.
- Dec. 21, 20:00-23:20 MCC CAMAC problem.
- Dec. 21, 23:20-23:50 Beam restoration and position check.
- Dec. 22, 08:00 beam off for holiday shut down.
- Total accummulated charge for pi+, Vertical pumping: 5.94 C.
Plans for the next few days:
- Beam studies: Monday 8-10 am
- Controlled Access and survey after beam studies
- Target polarimetry tests
- BigBite Shower calibrations
- Water calibration